# Discovery of a nova super-remnant surrounding the recurrent nova LMCN 1971-08a in the Large Magellanic Cloud

### v1

@@ 4.3 Confirmation of recurrent nature
LMCN 1971-08a is deemed a recurrent nova on account of its spatial coincidence with the location of nova N LMC 1971b (Bode et al., 2016). Though possibly some skepticism remained about the recurrent nature of the nova; for example, the entry for LMC 2009-02 (another identifier for LMCN 1971-08a) in Table 7 of Orio et al. (2021) is labelled as ‘perhaps’ a RN, unlike other accepted RNe in the LMC. Therefore, the discovery of a NSR around this nova further supports the status of an observationally-defined RN, but from a different approach.
In addition, owing to the existence of such a NSR, LMCN 1971-08a may have a much shorter recurrence period than the 38 years currently presumed and may therefore be observed in eruption much sooner than 2047.
@@ 5 Summary
In this Letter, we detailed the discovery of a nova super-remnant in the LMC hosted by the recurrent nova LMCN 1971-08a. The structure is spatially coincident with the location of the nova, is 200 pc in diameter making it the largest NSR found and is visible in narrowband H and emission indicative of shock-ionisation but with negligible emission. A corresponding structure is apparent in Hi data. Joining the growing collection of such NSRs, this NSR is the first to be discovered in the LMC, the brightest NSR to date and only the second extragalactic nova shell found after the M31 NSR (Darnley et al., 2019).
@@ A.1 NSR missed in previous work
The LMC was surveyed in previous work (Healy-Kalesh et al. 2024b) through a targeted search around the four recurrent novae using the Faulkes Telescope South but no NSRs were uncovered. While arcminute FOV images were provided in Healy-Kalesh et al. (2024b) around LMCN 1971-08a, the field of view of the Faulkes Telescope South facility used was arcminutes (chosen to search the surroundings of the RNe for this reason by Healy-Kalesh et al. (2024b)). As such, the region observed was large enough to capture a component of the NSR, however the relatively shorter exposure time of the observations (s for H) and therefore fainter emission, in addition to (and more importantly) only a component of the NSR being captured, indicates why this structure was overlooked in the previous work. The FOV of the Faulkes Telescope South observation from Healy-Kalesh et al. (2024b) is shown in Figure 6. Now with knowledge of the gross structure of the NSR, the southwest shell edge can be made out when comparing with the DeMCELS image in Figure 6, as can the structure to the west of the nova’s position.
@@ A.3 Comparison of line ratios with other possible structures
The ratios of emission lines can be used as diagnostic tools for helping to differentiate between different extended structures. As such, in Figure 7, we compare the ratio for the nova super-remnant with the same ratio for three LMC supernova remnants, J0450-7050 (a.k.a. Veliki; Smeaton et al. 2025) and J0534-7033 and J0506-6541 (from Zangrandi et al. 2024); the LMC Hii region, L77 DEM L45 (from Pellegrini et al. 2012); and the LMC planetary nebula, SMP LMC 80 (from Tosi et al. 2024), in order to differentiate the shell from other structures. We also provide similar comparative plots for the ratios and in Figure 8 and Figure 9, respectively.
@@
The ratios were calculated for a number of apertures (with radii of ) along a slit passing through the structure (illustrated in Figures 7, 8 and 9). As the structures are different sizes, the length of the line (and therefore the number of apertures used) were scaled accordingly such that the ratios can be compared. The slit shown in Figure 7 is long with 100 apertures for the nova super-remnant; long with 4 apertures for planetary nebula SMP LMC 80 (as PNe in the LMC are compact); long with 80 apertures for Hii region L77 DEM L45; long with 40 apertures for SNR J0450-7050; long with 25 apertures for SNR J0534-7033; and long with 50 apertures for SNR J0506-6541. In the ratio plots of Figures 7, 8 and 9, the axes have been normalised.
Supernova remnants:
@@
The emission line ratio for supernova remnants is typically high ( and see e.g. Yew et al. 2021; Caldwell et al. 2025) as a result of enhanced ionisation and excitation of sulphur from radiative shocks (Zangrandi et al. 2024). In addition to this, SNRs also exhibit strong emission as a result of shock heating (Yew et al. 2021). While the nova super-remnant does exhibit a ratio on par with the example supernova remnants ( and see Figure 8), there is a large difference in the ratios for the NSR and the SNRs, with the NSR emission being much fainter () within the NE and SW component of the shell (between and on the axis of Figure 7).
Planetary nebulae:
@@
Planetary nebulae have bright emission (Guerrero et al. 2013, and see the example PN in Figure 7) as a result of fast outflows shocking the surrounding low-density material (Guerrero et al. 2013) and weak emission (see, e.g. Caldwell et al. 2025). However, as can be seen in Figure 7, the nova super-remnant is faint in (with in the NE and SW component of the shell) and has moderate levels of emission (see Figure 8).
Hii regions:
@@
Typical Hii regions have low ratios as a result of the sulphur in such structures existing primarily as doubly-ionised sulphur (Yew et al. 2021) from the gas being photoionised (Zangrandi et al. 2024). As can be seen in Figure 8, the nova super-remnant has a higher ratio across the shell than the ratio of a typical Hii region. In addition, as illustrated with the example Hii region in Figure 7, Hii regions have high levels of due to photoionisation from hot central stars (Peimbert et al. 2017) however the discovered structure does not with for the NSR compared to for the Hii region.

### v2

@@ 4.3 Confirmation of recurrent nature
LMCN 1971-08a is deemed an RN on account of its spatial coincidence with the location of nova N LMC 1971b (Bode et al., 2016). Nevertheless, some scepticism has remained regarding its recurrent nature; for example, the entry for LMC 2009-02 (another identifier for LMCN 1971-08a) in Table 7 of Orio et al. (2021) is labelled as ”perhaps” an RN, unlike other confirmed RNe in the LMC. Thus, the discovery of an NSR around this nova further supports its classification as an observationally defined RN, albeit from a different approach. In addition, owing to the existence of such an NSR, LMCN 1971-08a may have a much shorter recurrence period than the 38 years currently presumed and may therefore be observed in eruption much sooner than 2047.
@@ 5 Summary
In this Letter, we detailed the discovery of a nova super-remnant in the LMC, hosted by the recurrent nova LMCN 1971-08a. The structure is spatially coincident with the nova’s location, spans 200 pc in diameter – making it the largest NSR found to date – and is visible in narrowband H and emission indicative of shock-ionisation, but exhibits negligible emission. A corresponding structure is apparent in Hi data. Adding to the growing collection of such NSRs, this is the first to be discovered in the LMC, the brightest NSR observed to date, and only the second extragalactic nova shell identified after the M31 NSR.
@@ A.1 NSR missed in previous work
The LMC was surveyed in previous work (Healy-Kalesh et al. 2024b) through a targeted search around the four recurrent novae using the Faulkes Telescope South but no NSRs were uncovered. While arcminute FOV images were provided in Healy-Kalesh et al. (2024b) around LMCN 1971-08a, the field of view of the Faulkes Telescope South facility used was arcminutes (chosen to search the surroundings of the RNe for this reason by Healy-Kalesh et al. 2024b). As such, the region observed was large enough to capture a component of the NSR, however the relatively shorter exposure time of the observations (s for H) and therefore fainter emission, in addition to (and more importantly) only a component of the NSR being captured, indicates why this structure was overlooked in the previous work. The FOV of the Faulkes Telescope South observation from Healy-Kalesh et al. (2024b) is shown in Fig. 6. Now with knowledge of the gross structure of the NSR, the southwest shell edge can be made out when comparing with the DeMCELS image in Fig. 6, as can the structure to the west of the nova’s position.
@@ A.3 Comparison of line ratios with other possible structures
The ratios of emission lines can be used as diagnostic tools for helping to differentiate between different extended structures. As such, in Fig. 7, we compare the ratio for the nova super-remnant with the same ratio for three LMC supernova remnants, J0450-7050 (a.k.a. Veliki; Smeaton et al. 2025) and J0534-7033 and J0506-6541 (from Zangrandi et al. 2024); the LMC Hii region, L77 DEM L45 (from Pellegrini et al. 2012); and the LMC planetary nebula, SMP LMC 80 (from Tosi et al. 2024), in order to differentiate the shell from other structures. We also provide similar comparative plots for the ratios and in Figs. 8 and 9, respectively.
@@
The ratios were calculated for a number of apertures (with radii of ) along a slit passing through the structure (illustrated in Figs. 7, 8, and 9). As the structures are different sizes, the length of the line (and therefore the number of apertures used) were scaled accordingly such that the ratios can be compared. The slit shown in Fig. 7 is long with 100 apertures for the nova super-remnant; long with 4 apertures for planetary nebula SMP LMC 80 (as PNe in the LMC are compact); long with 80 apertures for Hii region L77 DEM L45; long with 40 apertures for SNR J0450-7050; long with 25 apertures for SNR J0534-7033; and long with 50 apertures for SNR J0506-6541. In the ratio plots of Figs. 7, 8, and 9, the -axes have been normalised.
Supernova remnants
@@
The emission line ratio for supernova remnants is typically high ( and see, e.g., Yew et al. 2021; Caldwell et al. 2025) as a result of enhanced ionisation and excitation of sulphur from radiative shocks (Zangrandi et al. 2024). In addition to this, SNRs also exhibit strong emission as a result of shock heating (Yew et al. 2021). While the nova super-remnant does exhibit a ratio on par with the example supernova remnants ( and see Fig. 8), there is a large difference in the ratios for the NSR and the SNRs, with the NSR emission being much fainter () within the NE and SW component of the shell (between and on the axis of Fig. 7).
Planetary nebulae
@@
Planetary nebulae have bright emission (Guerrero et al. 2013, and see the example PN in Fig. 7) as a result of fast outflows shocking the surrounding low-density material (Guerrero et al. 2013) and weak emission (see, e.g., Caldwell et al. 2025). However, as can be seen in Fig. 7, the nova super-remnant is faint in (with in the NE and SW component of the shell) and has moderate levels of emission (see Fig. 8).
Hii regions
@@
Typical Hii regions have low ratios as a result of the sulphur in such structures existing primarily as doubly-ionised sulphur (Yew et al. 2021) from the gas being photoionised (Zangrandi et al. 2024). As can be seen in Fig. 8, the nova super-remnant has a higher ratio across the shell than the ratio of a typical Hii region. In addition, as illustrated with the example Hii region in Fig. 7, Hii regions have high levels of due to photoionisation from hot central stars (Peimbert et al. 2017) however the discovered structure does not with for the NSR compared to for the Hii region.